### v1

@@ 4.2 Defected graphene
Doping and defects play an important role in tuning the electronic properties of materials[38, 39, 40, 41, 42, 43]. Within the framework of first-principles calculations, such systems are usually modeled by supercells. Fig. 6(a) shows the geometric structure of defected graphene derived from a 5 $\times$ 5 supercell, for which the BZ is shown in Fig. 6(b). Fig. 6(c) shows the band structure for the defected 5 $\times$ 5 graphene. It looks completely different from the band structure of the perfect graphene. In addition, there are two new features. One is that there are dispersionless bands near the Fermi level. The other one is that there is no longer a gapless Dirac cone at K. Instead, it has a band gap. By projecting the wavefunctions of the defected graphene onto the $\mathbf{k}$-points in the first BZ of the corresponding primitive cell, the dispersionless bands are gone and a band gap in the Dirac cone at K can be seen. Moreover, there are minigaps in the electronic bands. This effect is due to the perturbation of the defects.
@@ 4.3 Graphene bilayer on 6H-SiC
Now we show an application of KPROJ in interfaces. We performed calculations for a graphene bilayer (graphene-2L) on 6H-SiC(0001), which has two types of terminations, i.e., Si and C terminations[44, 45, 46, 47, 48, 49, 50, 51]. Here, we consider an interface structure consisting of a 2 $\times$ 2 supercell of the graphene-2L on a $\sqrt{3}\times\sqrt{3}$ Si-terminated 6H-SiC(0001), which is saturated by hydrogen atoms. The geometric structure is shown in Fig. 7(a), for which the BZs are shown in Fig. 7(b). The calculated bands along the BZ of the graphene primitive cell are shown in Fig. 7(c). Note that there are two sets of band-crossing along the path from $\Gamma$-K. One is located at a $\mathbf{k}$-point (named $k_{q}$) in the middle of $\Gamma$-K and the other emerges at K. The extra band-crossing at $k_{q}$ is due to band folding, which still exists in the band structure weighted by the contribution of graphene shown in Fig .7(d). Then, we perform $k$-projection over the wavefunctions in the overlayer. The wavefunctions in the spatial window defined in Fig. 7(a) are chosen for the calculations. Then one can obtain an "effective" band structure for the bilayer graphene. This main profile of the unfolded band structure shows great similarity with that of the freestanding systems, except for minigaps in the electronic bands. Therefore, one can deduce that the Si-terminated 6H-SiC(0001) with H-saturation has a minor effect on the band structure of the overlayer.
@@ 4.4 Bulk projection of surfaces
Figure 8 shows the calculation for Ag(111). We build a slab with a thickness of 113 Å, which has 48 unit cells of Ag(111) along $z$ [Fig 8(a)]. We then remove two Ag(111) layers from each side of the slab to make a surface structure. By calculating the dispersions along $\Gamma-z$, one obtains dispersionless electronic bands [see fig. 8(d)]. By projecting the wavefunctions onto the $\mathbf{k}$-points in the BZ of the primitive cell of Ag bulk [Fig. 8(e)], one can obtain a band structure pretty much similar to that for the bulk phase. Note that the band crossing the Fermi level gets broadened as the energy goes to high energies, for which the surface effect plays a role. One can conveniently investigate the $k_{z}$ dependence of the band structure for the slab by projecting its wavefunctions onto the $k$-points in the BZ of the Ag bulk. The unfolded bands with different $k_{z}$s were shown in fig 8(f-i).
@@ 4.5 Phonons of defected graphene and moiré superlattices
Our KPROJ program also allows for unfolding phonon dispersions for doped materials and interfaces. We now show its application in the defected graphene and moiré superlattice with a twist angle of 9°. We use the small displacement method implemented in PHONOPY to obtain the phonon dispersions. The atomic forces induced by the displaced atom are calculated using LAMMPS[52] with a machine learning potential[53, 54, 55, 56, 57, 58], which was also used for the geometric relaxation. To facilitate comparison, fig. 9(a) shows the phonon spectrum for the graphene primitive cell. Figs. 9(b) and (c) show the folded and unfolded phonon spectrums for the defected graphene discussed in subsection 4.2. The unfolded phonon dispersions overall look very much similar to those for the free-standing graphene. However, one can also see many minigaps, which is due to the effects of carbon defects.
We also performed layer $k$-projection for graphene moiré superlattice, that is, we project the eigenvectors contributed by the top layer onto the $q$-point in the BZ of the graphene primitive cell. The results are shown in fig. 9(d). A comparison with fig. 9(c) reveals that the twisting and stacking have a much weaker effect than defects on the phonon dispersions of graphene.
@@ 6 Acknowledgement
This work was supported by the National Natural Science Foundation of China (Grant No. 12174098).

### v3

@@ 4.3 Defected graphene
Doping and defects play an important role in tuning the electronic properties of materials[46, 47, 48, 49, 50, 51]. Within the framework of first-principles calculations, such systems are usually modeled by supercells. Figure 10(a) shows the geometric structure of the defected graphene derived from a 5 $\times$ 5 supercell, for which the BZ is shown in Fig. 10(b). Figure 10(c) shows the band structure for the defected 5 $\times$ 5 graphene. Besides the folding, which makes it difficult to see the graphene bands, there are two new features: dispersionless bands near the Fermi level (as expected for a localized isolated defect), and the apparent disappearance of the Dirac cone. By projecting the wavefunction of the defected graphene onto the $k$-points in the first BZ of the corresponding primitive cell, (i) the bulk graphene states are recovered; (ii) there are two defect states, one is distributed over the whole BZ, and the other lower one is in the gap opened up in the lower part of the Dirac cone, and is mainly derived from those pure graphen states; and (iii) there are additional minigaps in the electronic bands due to defects.
@@ 4.4 Graphene bilayer on 6H-SiC
Now we show an application of KPROJ in interfaces. We performed calculations for a graphene bilayer (graphene-2L) on 6H-SiC(0001), which has two types of terminations, i.e., Si and C terminations[52, 53, 54, 55, 56, 57, 58, 59]. Here, we consider an interface structure consisting of a 2 $\times$ 2 supercell of the graphene-2L on a $\sqrt{3}\times\sqrt{3}$ Si-terminated 6H-SiC(0001), which is saturated by hydrogen atoms. The geometric structure is shown in Fig.11(a), for which the BZs are shown in Fig.11(b). The calculated bands along the BZ of the graphene primitive cell are shown in Fig.11(c). Note that there are two sets of band crossings along the path from $\Gamma$-K. One is located at a $k$-point (named $k_{q}$) in the middle of $\Gamma$-K and the other emerges at K. The extra band crossing at $k_{q}$ is due to band folding, which still exists in the band structure weighted by the contribution of graphene shown in Fig. 11(d). Then, we perform $k$-projection over the wave functions in the overlayer. The wave functions in the spatial window defined in Fig. 11(a) are chosen for the calculations. Then one can obtain an "effective" band structure for the bilayer graphene. The main profile of the unfolded band structure shows great similarity with that of the freestanding systems, except for minigaps in the electronic bands. Therefore, one can deduce that the Si-terminated 6H-SiC(0001) with H-saturation has a minor effect on the band structure of the overlayer.
@@ 4.5 Bulk projection of surfaces
Figure 12 shows the calculation for Ag(111). We build a slab with a thickness of 113 Å, which has 48 unit cells of Ag(111) along $z$ [Fig12(a)]. We then remove two Ag(111) layers from each side of the slab to make a surface structure. By calculating the dispersions along $\Gamma-Z$, one obtains dispersionless electronic bands [see Fig.12(d)]. By projecting the wave functions onto the $k$-points in the BZ of the primitive cell of bulk Ag [Fig. 12(e)], one obtains a band structure similar to that of the bulk. Note that the band crossing the Fermi level gets broadened as the energy goes to high energies, for which the surface effect plays a role. One can conveniently investigate the $k_{z}$ dependence of the band structure for the slab by projecting its wave functions onto the $k$-points in the BZ of the Ag bulk. The unfolded bands with different $k_{z}$s are shown in Fig.12(f-i).
@@ 4.6 Phonons of defected graphene and Moiré superlattices
Our KPROJ program also allows for unfolding phonon dispersions for doped materials and interfaces. We now show its application in the defected graphene and Moiré superlattice with a twist angle of 9°. We use the small displacement method implemented in PHONOPY to obtain the phonon dispersions. The atomic forces induced by the displaced atom are calculated using LAMMPS[60] with a machine learning potential[61, 62, 63, 64, 65, 66], which was also used for the geometric relaxation. To facilitate comparison, Fig. 13(a) shows the phonon spectrum for the graphene primitive cell, and Figs. 13(b) and (c) show the folded and unfolded phonon spectrums for the defected graphene, for which the geometric structure is shown in Fig. 10(a). The unfolded phonon dispersions look similar to those for free-standing graphene. However, one can also see many minigaps that are due to the effects of carbon defects.
We also performed layer $k$-projection for graphene Moiré superlattice, i.e., we project the eigenvectors contributed by the top layer onto the $q$-point in the BZ of the graphene primitive cell. The results are shown in Fig. 13(d). A comparison with Fig. 13(c) reveals that the twisting and stacking have a much weaker effect than do defects on the phonon dispersions of graphene.
@@ 6 Acknowledgement
This work was supported by the National Natural Science Foundation of China (No. 12174098 to M.C.) and the US National Science Foundation (No. DMREF 2323857 to M.W.).